OpenStudy (anonymous):

Is this correct ? Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x^2 - 6x + 2 = 0 A. {0.4, 5.6} B. {0.5, 4.8} C. {1.4, 2.4} D. no real number solution The answer is D.

OpenStudy (anonymous):

answer is \[3\pm\sqrt{7}\]

OpenStudy (anonymous):

so i was correct ?

OpenStudy (anonymous):

No the answer is a real number

OpenStudy (anonymous):

\[x^2-6x+2=0\] \[x^2-6x=-2\] \[(x-3)^2=-2+3^2=-2+9=7\] \[x-3=\pm\sqrt{7}\] \[x=3\pm\sqrt{7}\]

OpenStudy (anonymous):

so C ?

OpenStudy (anonymous):

calculator gives approximations as .35 and 5.6 so it is A